# Charting the Landscape of Nefarious Uses of Generative Artificial Intelligence for Online Election Interference

### v3

@@ Introduction
The advent of GenAI is revolutionizing fields from human-AI teaming to creative content generation, enabling new forms of human-computer interaction and augmenting human capabilities in unprecedented ways (Brown et al., 2020; OpenAI, 2023; Rahwan et al., 2019). However, these advancements come with significant risks: here we focus on the problem of interference with democracy (Bovet and Makse, 2019; Ozmen Garibay et al., 2023) within the broader context of AI ethics (Independent High-Level Expert Group on Artificial Intelligence, 2019; Mittelstadt et al., 2016). The primary objective of this study is to investigate the nefarious applications of GenAI in online election interference, exploring how these technologies can disrupt democratic processes. We aim to provide a comprehensive overview of the threats posed by GenAI, illustrating dual nature paradox that this technology poses, by means of a literature review and analysis to identify key trends and case studies (Barrett et al., 2023).
@@
Online election interference refers to the manipulation of digital information and communication technologies to influence the outcome of elections (Tenove et al., 2018; Kreps and Kriner, 2023). This can include activities such as spreading misinformation, creating fake news, and orchestrating social media campaigns to sway public opinion. The introduction of GenAI into this arena has magnified these threats by enabling more sophisticated and scalable forms of interference (Goldstein et al., 2023; Mozes et al., 2023; Shoaib et al., 2023). Generative AI, with its ability to create realistic text, images, and videos, presents new opportunities for malicious actors to deceive and manipulate the public. For example, deepfake technology can generate realistic videos of political figures saying or doing things they never did, eroding public trust in authentic information sources (see Fig. 1) (Seymour et al., 2023; Barrett et al., 2023; Kenthapadi et al., 2023; Lorenz et al., 2023; Gabriel et al., 2024). Moreover, the widespread use of AI-powered botnets can amplify misinformation and create the illusion of widespread support or opposition to certain political ideas, further distorting the democratic process (Unver, 2018; Yang and Menczer, 2023; Kreps and Kriner, 2023; Coeckelbergh, 2023).
@@
In our review, we identified several key case studies that illustrate the impact of GenAI on election interference (Yang et al., 2019; Kreps et al., 2022; Chennupati, 2024). For instance, during the 2016 U.S. presidential election, AI-generated botnets played a significant role in spreading misinformation and influencing public opinion (Bessi and Ferrara, 2016; Howard et al., 2018; Shao et al., 2018; Boichak et al., 2018). Similarly, deepfake videos have been used to create false narratives about political candidates, affecting voter perceptions and trust (Vaccari and Chadwick, 2020; Dobber et al., 2021; Diakopoulos and Johnson, 2021; Seymour et al., 2023). The deployment of GenAI in online election interference poses a unique challenge to democracies worldwide. Traditional methods of combating misinformation and digital manipulation are often insufficient to address the sophisticated techniques enabled by GenAI. This necessitates a comprehensive understanding of the various ways in which GenAI can be exploited for nefarious purposes and the development of new strategies to mitigate these risks (Xu et al., 2023; Gupta et al., 2023; Ziems et al., 2023).
@@
This perspective article aims to shed light on the darker applications of GenAI in the context of online election interference. By examining the various nefarious uses of these technologies, we hope to highlight the urgent need for regulatory oversight, technological solutions, public awareness, and collaborative efforts to safeguard the integrity of democratic processes. Through this exploration, we seek to provide a foundation for policymakers, technologists, and civil society to develop robust strategies to protect elections from AI-driven interference, ensuring that the benefits of GenAI are not overshadowed by its potential for harm (Rudin, 2019; Cao et al., 2023).
@@
GenAI and LLMs, celebrated for their ability to process natural language and generate multimodal content, offer numerous benefits, including enhancing creativity and solving complex problems (Brown et al., 2020; OpenAI, 2023; Shen et al., 2024; Yao et al., 2024). These technologies have enabled advancements in various fields, from natural language understanding and translation to creative industries such as art and music (Cao et al., 2023; Franceschelli and Musolesi, 2023). However, these same capabilities can be exploited for malicious purposes, such as creating deepfakes, synthetic identities, and targeted misinformation campaigns (Barrett et al., 2023; Mozes et al., 2023; Kang et al., 2024).
@@
On the positive side, GenAI can assist in drafting documents, writing code, generating realistic images and videos, and even conducting complex data analysis (Mandapuram et al., 2018; Bail, 2024; Li et al., 2024). These applications have immense potential to boost productivity and innovation across different sectors (Ooi et al., 2023; Kanbach et al., 2024). For example, AI-generated content can support marketing efforts by creating personalized advertisements or generating engaging social media posts. In healthcare, GenAI can help in diagnosing diseases by analyzing medical images or generating detailed patient reports (Davenport and Kalakota, 2019; Sai et al., 2024).
@@
Despite these benefits, the dark side of GenAI cannot be ignored (Wach et al., 2023; Kang et al., 2024). The same technology that generates realistic images and videos for legitimate purposes can also create deepfakes, which are manipulated media that convincingly alter appearances or actions of individuals. Deepfakes can be used to spread false information, damage reputations, and even blackmail individuals (Vaccari and Chadwick, 2020; Diakopoulos and Johnson, 2021; Seymour et al., 2023). The ease with which these deepfakes can be created poses a significant threat to the authenticity of information shared online. Similarly, LLMs can be used to generate convincing fake news articles, social media posts, and even entire personas (Chen and Shu, 2023; Augenstein et al., 2024; Choi and Ferrara, 2024a, b; Quelle and Bovet, 2024). These synthetic identities can be employed to infiltrate social networks, manipulate public opinion, and disrupt democratic processes (Haider et al., 2023; Luceri et al., 2024). AI-generated text can mimic human writing so effectively that distinguishing between genuine and fake content becomes challenging (Zellers et al., 2019; Augenstein et al., 2024; Quelle and Bovet, 2024; Jones and Bergen, 2024). Moreover, GenAI’s ability to tailor messages to specific demographics makes it a powerful tool for targeted misinformation campaigns. These campaigns can be designed to exploit societal divisions, amplify existing biases, and influence voter behavior (Boichak et al., 2018; Sap et al., 2019; Bovet and Makse, 2019; Goldstein et al., 2023; Ezzeddine et al., 2023; Hofmann et al., 2024). The targeted nature of these campaigns increases their effectiveness and the potential harm they can cause.
@@
The dual nature of GenAI highlights the need for a balanced approach in leveraging these technologies. While the benefits are substantial, the risks are equally significant (Weidinger et al., 2021, 2022; Barrett et al., 2023; Ferrara, 2024b). As we continue to develop and deploy GenAI, it is crucial to implement safeguards that prevent its misuse (Rudin, 2019). This includes developing robust detection mechanisms, establishing ethical guidelines, and promoting public awareness about the potential dangers of AI-generated content (Mittelstadt et al., 2016; Shahriari and Shahriari, 2017; Independent High-Level Expert Group on Artificial Intelligence, 2019; Jobin et al., 2019; Smuha, 2019; Yang et al., 2019; Qiu et al., 2023; Zhou et al., 2023; Sison et al., 2023). In conclusion, while GenAI holds great promise for enhancing various aspects of society, it also presents substantial risks. Understanding the dual nature of these technologies is essential for developing strategies that maximize their benefits while mitigating their potential harms (Seymour et al., 2023; Mozes et al., 2023).
@@
To sift through the growing scholarly work about the risks of GenAI, we carried out a systematic literature review and analysis. The literature review involved systematically identifying and analyzing scholarly articles, reports, and case studies related to the use of GenAI in online election interference. We focused on sources that provide empirical evidence and theoretical insights into the nefarious applications of these technologies. The analysis was conducted to identify key trends, influential works, and research gaps in the field. We analyzed recent journal and conference publication patterns, for a comprehensive understanding of the topic’s evolution.
@@
Nefarious Applications in Election Interference
The integration of GenAI into the sphere of online election interference introduces numerous sophisticated threats to the integrity of democratic processes. These applications leverage the advanced capabilities of AI to manipulate information and influence voter behavior in unprecedented ways. Here, we explore the primary nefarious applications of GenAI in election interference (Table 1).
@@
One of the most alarming applications of GenAI is the creation of deepfakes and other forms of synthetic media. Deepfakes use AI to create hyper-realistic videos and audio recordings of individuals, often portraying them saying or doing things they never did. Fabricated content can be used to spread false information, damage reputations, and manipulate public opinion (Seymour et al., 2023).
The potential impact of deepfakes on elections is significant. For instance, a deepfake video of a political candidate making controversial statements or engaging in unethical behavior could go viral, swaying public perception and potentially altering the outcome of an election. The speed at which such content can spread on social media platforms exacerbates the problem, making it difficult to counteract the false information before it influences a large audience (Vosoughi et al., 2018; Pinto et al., 2024).
@@
AI-powered botnets represent another potent tool for election interference. Botnets are networks of automated accounts that can post, like, share, and comment on social media platforms, creating the illusion of widespread support or opposition for certain viewpoints (Yang and Menczer, 2023). These bots can be programmed to spread misinformation, amplify divisive content, and suppress legitimate discourse, thereby distorting the digital public sphere (Ferrara et al., 2016).
The sophistication of modern AI allows these bots to mimic human behavior convincingly, making it challenging to detect and eliminate them (Ferrara, 2023b). For example, during the 2016 U.S. presidential election, bots were used extensively to influence online discussions and spread false information, contributing to the overall misinformation ecosystem (Ferrara et al., 2016; Ferrara, 2018). Pozzana and Ferrara (2020) provided insights into the dynamics between bot and human behaviors, illustrating the increasing sophistication of bot activities; Luceri et al. (2019b, a) showed how these behaviors evolve as a function of political ideology. The use of such technology in future elections poses a continuing threat to the integrity of democratic processes (Yang et al., 2022).
@@
GenAI’s ability to synthesize realistic and persuasive text enables the creation of highly effective targeted misinformation campaigns. These campaigns are designed to exploit existing societal divisions and biases, tailoring messages to specific demographic groups to maximize their impact (Vosoughi et al., 2018). By leveraging data from social media platforms and other online sources, malicious actors can create personalized propaganda that resonates deeply with targeted individuals, influencing their opinions, emotions, and behaviors.
For example, during an election, a targeted misinformation campaign might disseminate false information about a candidate’s policies or personal life, aiming to dissuade certain voter groups from supporting them. The personalized nature of these messages makes them particularly effective, as they can address the specific concerns and biases of the recipients (Allcott and Gentzkow, 2017). This form of manipulation undermines the democratic process by skewing the electorate’s perception based on falsehoods.
@@
GenAI can also be used to create convincing synthetic identities and fake accounts, which can be employed to infiltrate online communities, spread misinformation, and gather intelligence on political opponents (Ferrara, 2023b, 2024a). These synthetic identities can come with complete profiles, including photos, bios, and posting histories, making them difficult to distinguish from real users (Westerlund, 2019; Ezzeddine et al., 2023). The creation of fake accounts can facilitate a range of malicious activities. For instance, these accounts can be used to spread false narratives, engage in coordinated attacks against political figures, or even influence online polls. The use of synthetic identities to manipulate online discourse presents a significant challenge for social media platforms and requires sophisticated detection and countermeasures (Haider et al., 2023).
In summary, the nefarious applications of GenAI in election interference are diverse and highly sophisticated. From creating deepfakes to orchestrating botnets, these technologies offer malicious actors powerful tools to disrupt democratic processes and influence voter behavior. Addressing these threats requires a multifaceted approach that includes technological solutions, regulatory measures, and public awareness initiatives (Ferrara, 2024a).
@@
Election Country Techniques Used Actors Involved Outcome References 2016 UK Bots, social media campaigns, fake news Pro-Brexit groups, foreign actors Influenced public opinion, supported Brexit campaign Howard and Kollanyi (2016) 2016 USA Botnets, fake news, social media manipulation Russian state-sponsored groups Influenced public opinion, increased polarization Ferrara et al. (2016); Vosoughi et al. (2018) 2016 Philippines Troll accounts, fake news production Domestic political actors, misinformation agents Manipulated public opinion, supported political candidates Ong and Cabanes (2018); Haider et al. (2023) 2017 France Deepfakes, hacked emails, social media campaigns Russian state-sponsored groups, domestic actors Attempted to discredit candidates, disrupted public discourse Ferrara (2017); Pierri et al. (2023) 2017 Spain Bots, fake news, social media manipulation Pro-separatist groups, foreign actors Influenced public opinion, supported Catalan independence Stella et al. (2018) 2017 Germany Bots, trolls, misinformation Russian state-sponsored groups Influenced public opinion, attempted to disrupt political stability Bennhold and Barry (2017) 2018 USA Social media bots, fake news, targeted ads Foreign actors, domestic political groups Influenced voter behavior, increased polarization Deb et al. (2019) 2014/16 & 2018/22 Brazil Social media bots, fake news, WhatsApp campaigns Domestic political groups Polarized electorate, influenced election results Arnaudo (2017); Pacheco (2023) 2018 Italy Social media bots, misinformation campaigns Various political actors, misinformation agents Influenced public opinion, affected election outcomes Caldarelli et al. (2020) 2019 Indonesia Spear phishing, cyber attacks, social media manipulation Various political actors, cyber agents Influenced voter behavior, disrupted electoral process Tapsell (2020) 2019 India and Pakistan Fake social media accounts, coordinated misinformation Various political parties, independent actors Spread misinformation, manipulated voter sentiment Anand and Jaggi (2019) 2020 USA Deepfakes, AI-generated fake news, targeted ads Foreign actors, political groups Increased misinformation, voter suppression efforts Ferrara (2020); Ferrara et al. (2020) 2020 Taiwan Digital civic participation, misinformation campaigns Various political actors, misinformation agents Spread of misinformation, influence on voter perception and behavior Chang et al. (2021)
@@
The misuse of GenAI in election interference poses several significant societal risks. These risks extend beyond the immediate impact on elections (cf. Tables 2 and 3) and their outcomes and have profound implications for democratic institutions, social cohesion, and public trust.
@@
The widespread use of deepfakes and misinformation can severely erode public trust in media, institutions, and the democratic process itself. When voters are exposed to manipulated content that appears authentic, their ability to discern truth from falsehood is compromised. This erosion of trust can lead to cynicism and apathy among the electorate, as people become increasingly skeptical of all information sources (Seymour et al., 2023). The impact of this distrust is far-reaching. It undermines the credibility of legitimate news organizations and government institutions, making it easier for malicious actors to spread further misinformation (Baribi-Bartov et al., 2024). Once trust is broken, restoring it is complex, requiring significant efforts from involved stakeholders like media organizations, governments, and broader society (Ferrara, 2024a).
@@
Targeted misinformation campaigns exacerbate existing societal divisions by exploiting and amplifying biases and prejudices. By delivering tailored messages that resonate with specific demographic groups, these campaigns can deepen political and social polarization. For instance, misinformation targeting different ethnic or social groups can inflame tensions and lead to increased hostility between communities (Vosoughi et al., 2018). Such polarization weakens the social fabric, making it more challenging to achieve consensus on critical issues. This division can manifest in various forms, from increased partisanship in political discourse to violent confrontations between opposing groups. The long-term consequences of such division are detrimental to social stability and cohesion (Allcott and Gentzkow, 2017).
@@
The ability to manipulate public opinion on a large scale poses a direct threat to the foundations of democracy. Elections are meant to reflect the informed choices of the electorate. However, when these choices are influenced by false information, the legitimacy of the electoral process is compromised. This manipulation undermines the principle of free and fair elections, a cornerstone of democratic governance (Vosoughi et al., 2018). Badawy et al. (2019b) examined the demographics of individuals most susceptible to online political manipulation, highlighting the vulnerability of certain groups.
Furthermore, the perception that elections can be easily manipulated through AI-driven misinformation campaigns can lead to decreased voter turnout. If people believe their votes do not matter or that the process is rigged, they are less likely to participate in the democratic process. This disengagement weakens democratic institutions and can lead to governance that is less representative of the people’s will (Ferrara, 2023b).
@@
The deployment of GenAI technologies in ways that disproportionately affect certain groups can exacerbate existing social inequalities (Ferrara, 2023a). For example, misinformation campaigns might target marginalized communities with content designed to suppress their voter turnout or sow confusion about voting processes (Calo, 2017). Such tactics exploit the vulnerabilities of these communities, further entrenching their disenfranchisement. Moreover, access to sophisticated AI technologies is often concentrated among well-resourced entities, such as state actors or large corporations, giving them disproportionate power to influence public opinion and elections. This concentration of power can lead to a democratic deficit, where the interests of a few overshadow the voices of the many (Gillespie, 2018).

### v5

@@ Introduction
The advent of GenAI is revolutionizing fields from human-AI teaming to creative content generation, enabling new forms of human-computer interaction and augmenting human capabilities in unprecedented ways (gpt3, ; gpt4, ; rahwan2019machine, ). However, these advancements come with significant risks: here we focus on the problem of interference with democracy (bovet2019influence, ; ozmen2023six, ) within the broader context of AI ethics (EUAIethics, ; mittelstadt2016ethics, ). The primary objective of this study is to investigate the nefarious applications of GenAI in online election interference, exploring how these technologies can disrupt democratic processes. We aim to provide a comprehensive overview of the threats posed by GenAI, illustrating dual nature paradox that this technology poses, by means of a literature review and analysis to identify key trends and case studies (barrett2023identifying, ).
@@
Online election interference refers to the manipulation of digital information and communication technologies to influence the outcome of elections (tenove2018digital, ; kreps2023ai, ). This can include activities such as spreading misinformation, creating fake news, and orchestrating social media campaigns to sway public opinion. The introduction of GenAI into this arena has magnified these threats by enabling more sophisticated and scalable forms of interference (goldstein2023generative, ; mozes2023use, ; shoaib2023deepfakes, ). Generative AI, with its ability to create realistic text, images, and videos, presents new opportunities for malicious actors to deceive and manipulate the public. For example, deepfake technology can generate realistic videos of political figures saying or doing things they never did, eroding public trust in authentic information sources (see Fig. 1) (seymour2023beyond, ; barrett2023identifying, ; kenthapadi2023generative, ; lorenz2023initial, ; gabriel2024generative, ). Moreover, the widespread use of AI-powered botnets can amplify misinformation and create the illusion of widespread support or opposition to certain political ideas, further distorting the democratic process (unver2018artificial, ; yang2023anatomy, ; kreps2023ai, ; coeckelbergh2023democracy, ).
@@
In our review, we identified several key case studies that illustrate the impact of GenAI on election interference (yang2019arming, ; kreps2022all, ; chennupati2024threat, ). For instance, during the 2016 U.S. presidential election, AI-generated botnets played a significant role in spreading misinformation and influencing public opinion (bessi2016social, ; howard2018algorithms, ; shao2018spread, ; boichak2018automated, ). Similarly, deepfake videos have been used to create false narratives about political candidates, affecting voter perceptions and trust (vaccari2020deepfakes, ; dobber2021microtargeted, ; diakopoulos2021anticipating, ; seymour2023beyond, ). The deployment of GenAI in online election interference poses a unique challenge to democracies worldwide. Traditional methods of combating misinformation and digital manipulation are often insufficient to address the sophisticated techniques enabled by GenAI. This necessitates a comprehensive understanding of the various ways in which GenAI can be exploited for nefarious purposes and the development of new strategies to mitigate these risks (xu2023combating, ; gupta2023chatgpt, ; ziems2023can, ).
@@
This perspective article aims to shed light on the darker applications of GenAI in the context of online election interference. By examining the various nefarious uses of these technologies, we hope to highlight the urgent need for regulatory oversight, technological solutions, public awareness, and collaborative efforts to safeguard the integrity of democratic processes. Through this exploration, we seek to provide a foundation for policymakers, technologists, and civil society to develop robust strategies to protect elections from AI-driven interference, ensuring that the benefits of GenAI are not overshadowed by its potential for harm (rudin2019stop, ; cao2023comprehensive, ).
@@
GenAI and LLMs, celebrated for their ability to process natural language and generate multimodal content, offer numerous benefits, including enhancing creativity and solving complex problems (gpt3, ; gpt4, ; shen2024hugginggpt, ; yao2024tree, ). These technologies have enabled advancements in various fields, from natural language understanding and translation to creative industries such as art and music (cao2023comprehensive, ; franceschelli2023creativity, ). However, these same capabilities can be exploited for malicious purposes, such as creating deepfakes, synthetic identities, and targeted misinformation campaigns (barrett2023identifying, ; mozes2023use, ; kang2024exploiting, ).
@@
On the positive side, GenAI can assist in drafting documents, writing code, generating realistic images and videos, and even conducting complex data analysis (mandapuram2018investigating, ; bail2024can, ; li2024value, ). These applications have immense potential to boost productivity and innovation across different sectors (ooi2023potential, ; kanbach2024genai, ). For example, AI-generated content can support marketing efforts by creating personalized advertisements or generating engaging social media posts. In healthcare, GenAI can help in diagnosing diseases by analyzing medical images or generating detailed patient reports (davenport2019potential, ; sai2024generative, ).
@@
Despite these benefits, the dark side of GenAI cannot be ignored (wach2023dark, ; kang2024exploiting, ). The same technology that generates realistic images and videos for legitimate purposes can also create deepfakes, which are manipulated media that convincingly alter appearances or actions of individuals. Deepfakes can be used to spread false information, damage reputations, and even blackmail individuals (vaccari2020deepfakes, ; diakopoulos2021anticipating, ; seymour2023beyond, ). The ease with which these deepfakes can be created poses a significant threat to the authenticity of information shared online. Similarly, LLMs can be used to generate convincing fake news articles, social media posts, and even entire personas (chen2023combating, ; augenstein2024factuality, ; choi2024automated, ; choi2024fact, ; quelle2024perils, ). These synthetic identities can be employed to infiltrate social networks, manipulate public opinion, and disrupt democratic processes (haider2023detecting, ; luceri2024leveraging, ). AI-generated text can mimic human writing so effectively that distinguishing between genuine and fake content becomes challenging (zellers2019defending, ; augenstein2024factuality, ; quelle2024perils, ; jones2024people, ). Moreover, GenAI’s ability to tailor messages to specific demographics makes it a powerful tool for targeted misinformation campaigns. These campaigns can be designed to exploit societal divisions, amplify existing biases, and influence voter behavior (boichak2018automated, ; sap2019risk, ; bovet2019influence, ; goldstein2023generative, ; ezzeddine2023exposing, ; hofmann2024ai, ; ye2024auditing, ). The targeted nature of these campaigns increases their effectiveness and the potential harm they can cause.
@@
The dual nature of GenAI highlights the need for a balanced approach in leveraging these technologies. While the benefits are substantial, the risks are equally significant (weidinger2021ethical, ; weidinger2022taxonomy, ; barrett2023identifying, ; ferrara2024risks, ). As we continue to develop and deploy GenAI, it is crucial to implement safeguards that prevent its misuse (rudin2019stop, ). This includes developing robust detection mechanisms, establishing ethical guidelines, and promoting public awareness about the potential dangers of AI-generated content (mittelstadt2016ethics, ; shahriari2017ieee, ; EUAIethics, ; jobin2019global, ; smuha2019eu, ; yang2019arming, ; qiu2023latent, ; zhou2023ethical, ; sison2023chatgpt, ). In conclusion, while GenAI holds great promise for enhancing various aspects of society, it also presents substantial risks. Understanding the dual nature of these technologies is essential for developing strategies that maximize their benefits while mitigating their potential harms (seymour2023beyond, ; mozes2023use, ).
@@
To assess the growing body of scholarly work on the risks of generative AI (GenAI) in the context of online election interference, we conducted a systematic literature review. This review focused on identifying empirical studies, theoretical analyses, and case reports that examine the nefarious applications of GenAI technologies in political contexts. Our goal was to synthesize key trends, highlight influential contributions, and surface gaps in the existing research.
We searched major academic databases—Google Scholar, Scopus, and IEEE Xplore—using targeted keywords such as “generative AI,” “deepfakes,” “LLMs,” “election interference,” “misinformation,” and “synthetic media.” The review spanned publications from 2018 through early 2025 and included peer-reviewed articles, preprints, government and NGO reports, and selected white papers. To ensure quality and relevance, we included only sources that offered clear empirical findings or policy-oriented analysis. We excluded works focused solely on AI development without direct relevance to election interference. This structured approach allowed us to capture a diverse and representative picture of GenAI’s evolving threat landscape.
Nefarious Applications in Election Interference
The integration of GenAI into the sphere of online election interference introduces numerous sophisticated threats to the integrity of democratic processes. These applications leverage the advanced capabilities of AI to manipulate information and influence voter behavior in unprecedented ways (ferrara2020characterizing, ). Table 1 shows the primary nefarious applications of GenAI in election interference.
@@
Among the most alarming applications of generative AI (GenAI) is the creation of deepfakes and other synthetic media that convincingly mimic real individuals and events. Using technologies like generative adversarial networks (GANs), AI systems can fabricate highly realistic videos, audio recordings, and images—often depicting public figures saying or doing things they never actually did (seymour2023beyond, ). These false representations are increasingly deployed to spread disinformation, damage reputations, and manipulate political narratives.
The implications for democratic processes are profound. A single deepfake video showing a political candidate engaging in unethical behavior or making inflammatory remarks could go viral and influence voter perception before it can be effectively debunked. Social media platforms, with their speed and scale of dissemination, further exacerbate this challenge (vosoughi2018spread, ; pinto2024tracking, ). A notable real-world example involved a fabricated video of Ukrainian President Volodymyr Zelenskyy urging troops to surrender—swiftly proven false but widely circulated before corrections could reach many viewers.
Beyond video, synthetic audio poses a similarly insidious threat. Voice cloning tools can replicate tone, accent, and speech patterns, creating fake soundbites that are nearly indistinguishable from authentic recordings. Likewise, AI-generated images of fictitious people or events have been used to provoke public outrage or fabricate visual “evidence” for political claims.
It is crucial to distinguish these perceptual threats—rooted in visual and auditory realism—from text-based threats produced by large language models (LLMs). While LLMs can generate persuasive and scalable misinformation, synthetic media leverages sensory cues that typically elicit higher trust, making their influence especially potent. Text-based disinformation tends to be more subtle and diffuse, whereas synthetic media often delivers sharper, more visceral impacts.
Recognizing these modality-specific risks is essential for developing effective countermeasures. Tailored strategies must account for the unique properties of visual, audio, and text-based GenAI outputs—guiding both the design of detection technologies and the formulation of policy responses.
@@
AI-powered botnets represent another potent tool for election interference. Botnets are networks of automated accounts that can post, like, share, and comment on social media platforms, creating the illusion of widespread support or opposition for certain viewpoints (yang2023anatomy, ). These bots can be programmed to spread misinformation, amplify divisive content, and suppress legitimate discourse, thereby distorting the digital public sphere (ferrara2016rise, ).
The sophistication of modern AI allows these bots to mimic human behavior convincingly, making it challenging to detect and eliminate them (ferrara2023social, ). For example, during the 2016 U.S. presidential election, bots were used extensively to influence online discussions and spread false information, contributing to the overall misinformation ecosystem (ferrara2016rise, ; ferrara2018measuring, ). (pozzana2020measuring, ) provided insights into the dynamics between bot and human behaviors, illustrating the increasing sophistication of bot activities; (luceri2019evolution, ; luceri2019red, ) showed how these behaviors evolve as a function of political ideology. The use of such technology in future elections poses a continuing threat to the integrity of democratic processes (yang2022botometer, ).
@@
GenAI’s ability to synthesize realistic and persuasive text enables the creation of highly effective targeted misinformation campaigns. These campaigns are designed to exploit existing societal divisions and biases, tailoring messages to specific demographic groups to maximize their impact (vosoughi2018spread, ). By leveraging data from social media platforms and other online sources, malicious actors can create personalized propaganda that resonates deeply with targeted individuals, influencing their opinions, emotions, and behaviors (blas2024unearthing, ).
For example, during an election, a targeted misinformation campaign might disseminate false information about a candidate’s policies or personal life, aiming to dissuade certain voter groups from supporting them. The personalized nature of these messages makes them particularly effective, as they can address the specific concerns and biases of the recipients (allcott2017social, ). This form of manipulation undermines the democratic process by skewing the electorate’s perception based on falsehoods.
@@
GenAI can also be used to create convincing synthetic identities and fake accounts, which can be employed to infiltrate online communities, spread misinformation, and gather intelligence on political opponents (ferrara2023social, ; ferrara2024genai, ). These synthetic identities can come with complete profiles, including photos, bios, and posting histories, making them difficult to distinguish from real users (westerlund2019emergence, ; ezzeddine2023exposing, ). The creation of fake accounts can facilitate a range of malicious activities. For instance, these accounts can be used to spread false narratives, engage in coordinated attacks against political figures, or even influence online polls (memo4, ; minici2024iohunter, ; cinus2024exposing, ). The use of synthetic identities to manipulate online discourse presents a significant challenge for social media platforms and requires sophisticated detection and countermeasures (haider2023detecting, ).
In summary, the nefarious applications of GenAI in election interference are diverse and highly sophisticated. From creating deepfakes to orchestrating botnets, these technologies offer malicious actors powerful tools to disrupt democratic processes and influence voter behavior (memo6, ). Addressing these threats requires a multifaceted approach that includes technological solutions, regulatory measures, and public awareness initiatives (ferrara2024genai, ).
@@
Election Country Techniques Used Actors Involved Outcome References 2016 UK Bots, social media campaigns, fake news Pro-Brexit groups, foreign actors Influenced public opinion, supported Brexit campaign (howard2016bots, ) 2016 USA Botnets, fake news, social media manipulation Russian state-sponsored groups Influenced public opinion, increased polarization (ferrara2016rise, ; vosoughi2018spread, ) 2016 Philippines Troll accounts, fake news production Domestic political actors, misinformation agents Manipulated public opinion, supported political candidates (ong2018architects, ; haider2023detecting, ) 2017 France Deepfakes, hacked emails, social media campaigns Russian state-sponsored groups, domestic actors Attempted to discredit candidates, disrupted public discourse (ferrara2017disinformation, ; pierri2023propaganda, ) 2017 Spain Bots, fake news, social media manipulation Pro-separatist groups, foreign actors Influenced public opinion, supported Catalan independence (stella2018bots, ) 2017 Germany Bots, trolls, misinformation Russian state-sponsored groups Influenced public opinion, attempted to disrupt political stability (bennhold2017russia, ) 2018 USA Social media bots, fake news, targeted ads Foreign actors, domestic political groups Influenced voter behavior, increased polarization (deb2019perils, ) 2014/16 & 2018/22 Brazil Social media bots, fake news, WhatsApp campaigns Domestic political groups Polarized electorate, influenced election results (arnaudo2017computational, ; pacheco2023bots, ) 2018 Italy Social media bots, misinformation campaigns Various political actors, misinformation agents Influenced public opinion, affected election outcomes (caldarelli2020role, ) 2019 Indonesia Spear phishing, cyber attacks, social media manipulation Various political actors, cyber agents Influenced voter behavior, disrupted electoral process (tapsell2020weaponization, ) 2019 India and Pakistan Fake social media accounts, coordinated misinformation Various political parties, independent actors Spread misinformation, manipulated voter sentiment (anand2019whatsapp, ) 2020 USA Deepfakes, AI-generated fake news, targeted ads Foreign actors, political groups Increased misinformation, voter suppression efforts (ferrara2020bots, ; ferrara2020characterizing, ) 2020 Taiwan Digital civic participation, misinformation campaigns Various political actors, misinformation agents Spread of misinformation, influence on voter perception and behavior (chang2021digital, )
@@
Election interference poses several significant societal risks. These risks extend beyond the immediate impact on elections (see Table 2) and their outcomes and have profound implications for democratic institutions, social cohesion, and public trust (ratkiewicz2011detecting, ; broniatowski2018weaponized, ; stella2018bots, ; badawy2019characterizing, ; im2020still, ; chang2021digital, ).
In Tables 3 and 4 we present a typology-based overview of GenAI threats—organized by modality, target, objective, and use case, respectively for democratic and authoritarian contexts.
@@
The widespread use of deepfakes and misinformation can severely erode public trust in media, institutions, and the democratic process itself. When voters are exposed to manipulated content that appears authentic, their ability to discern truth from falsehood is compromised. This erosion of trust can lead to cynicism and apathy among the electorate, as people become increasingly skeptical of all information sources (seymour2023beyond, ). The impact of this distrust is far-reaching. It undermines the credibility of legitimate news organizations and government institutions, making it easier for malicious actors to spread further misinformation (baribi2024supersharers, ). Once trust is broken, restoring it is complex, requiring significant efforts from involved stakeholders like media organizations, governments, and broader society (ferrara2024genai, ).
@@
Targeted misinformation campaigns exacerbate existing societal divisions by exploiting and amplifying biases and prejudices. By delivering tailored messages that resonate with specific demographic groups, these campaigns can deepen political and social polarization. For instance, misinformation targeting different ethnic or social groups can inflame tensions and lead to increased hostility between communities (vosoughi2018spread, ). Such polarization weakens the social fabric, making it more challenging to achieve consensus on critical issues. This division can manifest in various forms, from increased partisanship in political discourse to violent confrontations between opposing groups. The long-term consequences of such division are detrimental to social stability and cohesion (allcott2017social, ).
@@
The ability to manipulate public opinion on a large scale poses a direct threat to the foundations of democracy. Elections are meant to reflect the informed choices of the electorate. However, when these choices are influenced by false information, the legitimacy of the electoral process is compromised (luceri2024susceptibility, ). This manipulation undermines the principle of free and fair elections, a cornerstone of democratic governance (vosoughi2018spread, ). (badawy2019falls, ) examined the demographics of individuals most susceptible to online political manipulation, highlighting the vulnerability of certain groups.
Furthermore, the perception that elections can be easily manipulated through AI-driven misinformation campaigns can lead to decreased voter turnout. If people believe their votes do not matter or that the process is rigged, they are less likely to participate in the democratic process. This disengagement weakens democratic institutions and can lead to governance that is less representative of the people’s will (ferrara2023social, ).
@@
The deployment of GenAI technologies in ways that disproportionately affect certain groups can exacerbate existing social inequalities (ferrara2023should, ). For example, misinformation campaigns might target marginalized communities with content designed to suppress their voter turnout or sow confusion about voting processes (calo2017artificial, ). Such tactics exploit the vulnerabilities of these communities, further entrenching their disenfranchisement. Moreover, access to sophisticated AI technologies is often concentrated among well-resourced entities, such as state actors or large corporations, giving them disproportionate power to influence public opinion and elections (pinto2024tracking, ). This concentration of power can lead to a democratic deficit, where the interests of a few overshadow the voices of the many (gillespie2018custodians, ).